# Continual Learning in Medical Imaging from Theory to Practice: A Survey and Practical Analysis

### v1

@@ 4.1 Trends
Consequently, most of the research has concentrated on regularization approaches to address continual learning challenges. Over half of the reviewed works opted for regularization-based methods to tackle this issue effectively. Furthermore, there has been a recent surge in dynamic model approaches, which, although are less memory efficient and complex to train, offer a simpler solution (van de Ven et al., 2022). Dynamic models, which involve maintaining an ensemble of models for various tasks, offer a straightforward means to handle the problem, contrasting with alternative approaches that may entail intricate learning processes.
@@ 4.2 Practicality of CL
4.2.1 Standardization
The need for continual learning in the medical domain is paramount, mainly because of data scarcity and privacy concerns. However, current research in this domain lacks a clear structure and standard methodology for conducting experiments. This eventually makes it difficult to compare different studies, which may not provide a complete analysis of the proposed methods. Moreover, the absence of standardized benchmarks further complicates the validation of CL algorithms in medical imaging. Consequently, there is a pressing need to establish common evaluation protocols and datasets to ensure fair comparison and robust assessment of CL techniques in real-world healthcare settings. Furthermore, in the medical domain, the scarcity of publicly available datasets is a significant obstacle. The primary dataset used in recent studies is the MedMNIST dataset (Yang et al., 2021). This dataset lacks a specific task structure, and different authors often choose different tasks and settings to evaluate their methods. Notably, even a small shift in the order of the tasks can cause a significant difference in overall performance. While most of the reviewed studies mention how they partitioned the datasets they used, the comparison becomes challenging owing to differences in settings, leading to unfair comparisons. Therefore, there is a clear need to develop standard datasets, comparison settings, and metrics to evaluate and compare the CL methods.
4.2.2 Task Boundaries
@@
4.2.3 Reproducibility and Computational Constraints
Reproducibility and making the source code public are highly important in CL in the medical domain. This ensured that the findings and models could be validated and used by other researchers to foster trust and confidence in the results. The open-source development of these methods is crucial for the advancement of CL as a field. Moreover, the computational aspect is important for scalability and efficiency of systems for larger populations. Unlike the natural dataset domain, medical domain models must be deployed on-site in the hospital so that they can access the local clinical data repository without any privacy risk. The computational resources at hospitals are very limited compared with the large cloud-based solutions used in other domains, further increasing the need for efficient solutions. Therefore, reproducibility and computational efficiency are essential for advancing CL in the medical domain and ultimately improving patient care and outcomes.
4.2.4 Regulatory Framework for Model Updates
Several regulations govern update protocols for deployed models (Moskalenko and Kharchenko, 2024). These regulations demand transparency in the update process, while ensuring that the model’s performance does not deteriorate significantly on previously learned tasks. In addition to transparency and performance preservation, compliance with regulatory requirements often requires addressing various aspects. Ensuring data privacy is paramount, with updates required to strictly adhere to data privacy regulations to safeguard sensitive information. Ethical considerations, including fairness, bias, and accountability, must also be considered when updating the models to maintain ethical standards. Moreover, rigorous validation procedures are essential to assess the impact of updates on model performance and to ensure compliance with regulatory standards.
5 Recommendation

### v2

@@ IV-A Trends
Consequently, most of the research has concentrated on regularization approaches to address continual learning challenges. Over half of the reviewed works opted for regularization-based methods to tackle this issue effectively. Furthermore, there has been a recent surge in dynamic model approaches, which, although are less memory efficient and complex to train, offer a simpler solution [18]. Dynamic models, which involve maintaining an ensemble of models for various tasks, offer a straightforward means to handle the problem, contrasting with alternative approaches that may entail intricate learning processes.
@@ IV-B Practicality of CL
IV-B1 Standardization
The need for continual learning in the medical domain is paramount, mainly because of data scarcity and privacy concerns. However, current research in this domain lacks a clear structure and standard methodology for conducting experiments. This eventually makes it difficult to compare different studies, which may not provide a complete analysis of the proposed methods. Moreover, the absence of standardized benchmarks further complicates the validation of CL algorithms in medical imaging. Consequently, there is a pressing need to establish common evaluation protocols and datasets to ensure fair comparison and robust assessment of CL techniques in real-world healthcare settings. Furthermore, in the medical domain, the scarcity of publicly available datasets is a significant obstacle. The primary dataset used in recent studies is the MedMNIST dataset [95]. This dataset lacks a specific task structure, and different authors often choose different tasks and settings to evaluate their methods. Notably, even a small shift in the order of the tasks can cause a significant difference in overall performance. While most of the reviewed studies mention how they partitioned the datasets they used, the comparison becomes challenging owing to differences in settings, leading to unfair comparisons. Moreover, the size of the images in this dataset makes it debatable whether it’s a benchmark that would transfer to clinical usability. Therefore, there is a clear need to develop standard datasets, comparison settings, and metrics to evaluate and compare the CL methods.
IV-B2 Task Boundaries
@@
IV-B3 Reproducibility and Computational Constraints
Reproducibility and making the source code public are highly important in CL in the medical domain. This ensures that the findings and models could be validated and used by other researchers to foster trust and confidence in the results. The open-source development of these methods is crucial for the advancement of CL as a field. Moreover, the computational aspect is important for scalability and efficiency of systems for larger populations. Unlike the natural dataset domain, medical domain models must be deployed on-site in the hospital so that they can access the local clinical data repository without any privacy risk. The computational resources at hospitals are very limited compared with the large cloud-based solutions used in other domains, further increasing the need for efficient solutions. Therefore, reproducibility and computational efficiency are essential for advancing CL in the medical domain and ultimately improving patient care and outcomes.
IV-B4 Regulatory Framework for Model Updates
Several regulations govern update protocols for deployed models [96]. These regulations demand transparency in the update process, while ensuring that the model’s performance does not deteriorate significantly on previously learned tasks. In addition to transparency and performance preservation, compliance with regulatory requirements often requires addressing various aspects. Ensuring data privacy is paramount, with updates required to strictly adhere to data privacy regulations to safeguard sensitive information. Ethical considerations, including fairness, bias, and accountability, must also be considered when updating the models to maintain ethical standards. Moreover, rigorous validation procedures are essential to assess the impact of updates on model performance and to ensure compliance with regulatory standards.
V Recommendation